A support cable on a bridge has an area of cross-section of 0.0085 m2
and a length of
35 m. It is made of high tensile steel whose Young’s modulus is 2.8 1011 Pa. The
tension in the cable is 720 kN. Calculate
a. the extension of the cable.
b. the strain energy stored in the cable.
1
Expert's answer
2021-08-26T19:04:04-0400

a)

"e=\frac{T}{EA}l\\e=\frac{720\cdot10^{3}}{(0.0085)(2.8\cdot10^{11})}35=0.011\ m"

b)


"E_e=0.5(720\cdot10^3)(0.011)=4000\ J"